# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 5.5 Truthfulness
One prominent drawback of LLMs is their tendency to confidently produce false statements. These statements fall into two main categories: 1) statements that contradict learned knowledge within models, a problem often related to models’ honesty; 2) statements that are factually incorrect and appear to be fabricated by models, a phenomenon commonly referred to as hallucination. In the following, we delve into various approaches that aim to understand aforementioned two behaviors by leveraging explainability tools.
5.5.1 Honesty
@@
Honesty of LLMs describes models’ ability to produce true statements based on their learned information, where dishonest behaviors significantly compromise the perceived trustworthiness of LLMs. Numerous studies have been undertaken to understand how and why dishonest behaviors happen by looking inside LLMs. One notable work attempts to distinguish dishonesty by training a classifier to predict the accuracy of statements (Azaria & Mitchell, 2023). The classifier is simply trained on top of activations from the hidden layers of LLMs. These activations are generated from true and false statements. The classifier reaches an accuracy range between 60% and 80%, which suggests that LLMs may be internally aware of the truthfulness of their outputs (Azaria & Mitchell, 2023). Furthermore, research by Campbell et al. (2023) localizes dishonesty behaviors at the level of attention heads. This study employs specially crafted prompts to induce dishonest responses, and then trains logistic classifiers on models’ activations over true/false statements. It also employs activation patching to substitute lying activations with honesty ones. Intriguingly, both approaches have witnessed the importance of layer 23--29 in flipping dishonesty behaviors. Besides, another popular method tries to study the geometric structure of true/false statements(Marks & Tegmark, 2023). Typically, these structures are visualized by projecting representations of statements onto two principal components. A clear linear structure and the truth directions can be derived to mitigate the dishonest behaviors.
5.5.2 Hallucinations
@@
Hallucinations in LLMs can arise due to poor data quality, biases, outdated knowledge, and the lack of explicit knowledge (Xu et al., 2024; Zhu et al., 2023b). However, whether LLMs are aware of their hallucination behaviors remains an open question. Recent work investigates this question by examining models’ hidden representation space (Duan et al., 2024). It examines three hidden states involving a question, its correct answer and its incorrect answer, which are used to compute an “awareness” score. This metric quantifies the uncertainty of LLMs regarding their own answers, finding that adversarially induced hallucination can increase models’ awareness. Additionally, Li et al. (2024b) illustrates the major differences between models’ output and their inner activations, identifying these discrepancies as a potential source of hallucination. By training linear probing classifiers on each attention head’s activations, the most specialized attention head is identified. An orthogonal probe is subsequently trained with the identified specialized probe. The experiments reveal that “truth” might exist in a subspace instead of a single direction (Li et al., 2024b). Another work investigates the source of hallucination by analyzing patterns of source token contributions through perturbations (Xu et al., 2023c). Their findings suggest that hallucinations may stem from the models’ excessive dependence on a restricted set of source tokens. Besides, the static distribution of source token contribution, termed as “source contribution staticity”, can be used as another indicator of hallucinations.
Building on the above insights into LLM hallucinations, Duan et al. (2024) apply PCA to derive the direction of the correct answer’s final hidden state, and enhance the hidden representations with this direction to reduce hallucinations. In contrast, Li et al. (2024b) adopts a different approach, by intervening on top-$K$ specialized attention heads, while minimizing the influence of the rest attention heads within models. Different from PCA that identifies a single principle direction, this work adopts two distinct techniques to find multiple directions of intervention. First, they use orthogonal vectors of each probe’s hyperplane, which is similar to PCA. Second, they leverage vectors that connect the mean of the true and false distributions (Li et al., 2024b). The vectors derived from mean shift has been demonstrated more effective than those from probe classifiers, which presents another feasible strategy for identifying directions of truth.
5.6 Challenges
@@
We discuss the challenges in employing explanations to improve models’ trustworthiness and enhance alignment from two aspects: 1) limitations of existing detection techniques, and 2) shortcomings of mitigation strategies based on explanations.
5.6.1 Challenges of Existing Detection Methods
Current detection methods primarily focus on the level of LLM modules, such as layers, attention heads, and representations. However, we still lack a finer-grained understanding of how knowledge is encoded within LLMs, such as at the level of individual neurons and weight parameters. Furthermore, we lack general and robust strategies to identify these knowledge reliably. For instance, to identify gender biases, attention heads might be examined and then the related heads pruned (Li et al., 2024b). This approach requires analyzing each model individually, rather than adopting a general approach. Moreover, existing localization approaches rely either on probing classifiers or casual scrubbing, which might not be reliable. For probing classifiers, the pre-designed biases used to train these classifiers are crucial to their performance. On the other hand, casual cleaning usually introduces new variables that complicate the analysis.
5.6.2 Challenges of Mitigation Strategies
@@
Since LLMs are trained on vast corpora of data, it is impractical to mitigate the aforementioned trustworthiness issues from the data perspective. It is widely believed that XAI, i.e. understanding the inner mechanisms of LLMs, is a fundamental way to solve these issues. Thus, mitigation methods for LLMs are typically developed based on explanations. Existing explanations are implemented using techniques from mechanistic interpretability and representation engineering (Zhao et al., 2024). While both streams of methods can alleviate these issues, they fail to fully address them. For example, principal component analysis (PCA) is often used to find geometric structures related to these issues, but cannot fix these issues. Another popular mitigation method is casual scrubbing, which introduces certain "positive" activations to replace those demonstrated to be responsible for specific issues. However, the identified directions and patched activations can only mitigate issues to a certain extent. Moreover, the changes to either representations or activations could also influence other aspects of models’ capabilities, which we are yet unable to evaluate.
6 LLM Enhancement via Explainable Prompting
@@
A key distinction between LLMs and traditional machine learning models lies in the LLMs’ ability to accept flexibly manipulated input data, namely prompts, during model inference (Liu et al., 2023a). LLMs generally give precedence to the information presented in these prompts when generating outputs. Therefore, to mitigate the opacity issue in LLM predictions, we can enhance prompts with understandable content, which is then prioritized over the LLMs’ inherent and implicit knowledge. These enriched prompts can include domain-specific insights, contextual information, or a step-by-step reasoning chain. In response, LLMs might reveal their decision-making processes during inference, which improves the explainability of their behaviors.
6.1 Chain of Thoughts (CoT) Prompting
The Chain of Thought (CoT) approach significantly enhances LLMs in tackling complex tasks (Wei et al., 2022). While LLMs are adept at generating human-like responses, they often lack transparency in their reasoning processes. This limitation makes it difficult for users to assess the credibility of the responses, especially for questions that require detailed reasoning.
@@ 6.1 Chain of Thoughts (CoT) Prompting
To bridge this gap, recent efforts have incorporated in-context learning with human-crafted explanations directly into prompts (Wei et al., 2022; Huang et al., 2023a; Yao et al., 2023b; Besta et al., 2023). Among these approaches, the CoT method stands out by employing explicit knowledge to guide the reasoning process. Formally, we define the language model as $f_{\theta}$, and input prompt as $X=\{x_{1},y_{1},x_{2},y_{2},...,x_{n}\}$, where $x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1}$ denote the example question-response pairs for in-context learning, and $x_{n}$ is the actual question. In a standard question-answering scenario, we have the model output as $y_{n}=\operatorname*{arg\,max}_{Y}p_{\theta}(Y|x_{1},y_{1},x_{2},y_{2},...,x_{n})$. This approach, however, does not provide insights into the reasoning process behind the answer $y_{n}$. Therefore, the CoT method proposes to include human-crafted explanations $e_{i}$ for the $i$-th in-context example, resulting in a modified input format $X=\{x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}\}$. Given the input, the model will output not only $y_{n}$ but also the generated explanation $e_{n}$:
$e_{n},y_{n}=\operatorname*{arg\,max}_{Y}p_{\theta}(Y|x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}).$ (5)
Besides allowing for a more transparent and understandable interaction with LLMs, the CoT approach is also practically useful as it augments LLMs’ functionality by opening a window for users to control the models’ thought processes. Specifically, the usefulness of CoT methods lies in several key aspects:
@@
  • Reducing Errors in Reasoning: By breaking down complex problems into a series of smaller tasks, CoT reduces errors in complex calculations and logic-oriented tasks, leading to a more precise resolution of intricate problems (Wei et al., 2022; Qin et al., 2023; Zhang et al., 2023; Wang & Zhou, 2024b).
@@
  • Facilitating Knowledge Distillation: The step-by-step reasoning processes derived from larger LLMs can serve as a specialized fine-tuning dataset for smaller LLMs. It allows smaller models to learn complex problem-solving by following explanations, effectively teaching them to tackle intricate questions with enhanced reasoning capabilities (Magister et al., 2022).
6.2 Extended Methods of Explainable Prompting
Advanced techniques beyond CoT methods have been developed to broaden the range of reasoning paths available to LLMs towards enhancing the transparency and understandability of the decision-making process (Yao et al., 2023b; Besta et al., 2023; Yao et al., 2023a; Dhuliawala et al., 2023; Lyu et al., 2023). We introduce several notable examples below.
Tree-of-Thoughts (ToT). Proposed by Yao et al. (2023b), ToT advances beyond the traditional linear Chain of Thought reasoning, offering a more versatile structure that allows models to navigate through multiple reasoning paths. ToT makes the reasoning process of LLMs more interpretable by closely aligning it with human thought processes, as humans naturally consider multiple options and possible outcomes in both forward planning and retrospective analysis to reach conclusions (Sloman, 1996; Stanovich, 1999). This capability enhances the capacity of LLMs to tackle complex challenges that require the ability to consider and reevaluate different strategies, such as devising game strategies or generating creative content. By simulating the way humans think and make decisions, ToT not only makes their thought process more understandable to human users, but also improves the models’ effectiveness in handling complex tasks.
Graph of Thoughts (GoT). Proposed by Besta et al. (2023), GoT transforms the output of LLMs into a graph format. This format visualizes information pieces as nodes and their connections as edges, enabling a more intricate and connected form of reasoning compared to previous methods CoT and ToT. By organizing data into nodes (individual concepts or pieces of information) and edges (relationship between these concepts), GoT makes the logical connections within complex systems more understandable (Yao et al., 2023a). This graphical representation brings several benefits for understanding complex information. Firstly, it enables dynamic modification of relationships between concepts, offering a clear visualization of how changing one element affects the others. This is crucial in fields like legal reasoning (Cui et al., 2023; Boche et al., 2024), scientific research (Ding et al., 2023; Choudhury et al., 2023), and policy analysis (Chen et al., 2023c), where the inter-dependencies between various factors can be intricate and subtle. Secondly, GoT enables an assessment of the significance of each node within the graph, providing insights into which pieces of information are most critical to the task. This level of adaptability and clarity makes GoT exceptionally powerful for analyzing and navigating complex information networks.
6.3 Case Study: Is CoT Really Making LLM Inferences Explainable?
6.3.1 Background and Experimental Settings
@@
Despite the apparent intuitiveness of the CoT prompt design, a critical question remains unanswered: Does CoT really make LLM inferences explainable? In other words, can the information provided through CoT faithfully reflect the underlying generation process of LLMs? We use multi-hop question-answering (QA) as the scenario to investigate this problem.
In QA systems, answering multi-hop questions remains a significant challenge. Instead of leveraging a single information source, multi-hop questions require synthesizing information from multiple pieces or sources of data into a coherent and logical sequence. While LLMs show good performance in single-hop QA tasks (Radford et al., 2019), their efficacy significantly declines in multi-hop situations (Tan et al., 2023; Kim et al., 2023a; Zhong et al., 2023). This discrepancy highlights the need for more advanced methods to effectively handle the intricacy of multi-hop reasoning.
@@
CoT Prompts for Multi-hop QA. To address the above challenge, our case study applies the CoT technique. CoT relies on high-quality templates as the in-context learning prompt, where we give an example as below. Here, $[x]$ denotes the test question. The “Thoughts” following each “Question” are step-by-step problem-solving statements for the multi-hop questions. The thoughts in the templates align the generation process of LLMs with human cognitive problem-solving patterns.
Question: What is the capital of the country where Plainfield Town Hall is located?
Thoughts: Plainfield Town Hall is located in the country of the United States of America. The capital of the United States is Washington, D.C.
Answer: Washington, D.C.
Question: Who has ownership of the developer of the Chevrolet Corvette (C4)?
Thoughts: The developer of Chevrolet Corvette (C4) is Chevrolet. Chevrolet is owned by General Motors.
Answer: General Motors
Question: [x]
@@
CoT Faithfulness for Explanation: To quantitatively measure the faithfulness of CoTs, we select fidelity as the corresponding metrics (Zhao et al., 2023b; Wachter et al., 2017):
@@
$Fidelity=\frac{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)-\mathbbm{1}\left(\hat{\mathbf{y}}_{i}^{mislead}=\mathbf{y}_{i}\right)\right)}{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)\right)}\times 100\%,$ (6)
@@
where $\mathbf{y}_{i}$ denotes the ground truth label, $\hat{\mathbf{y}}_{i}$ denotes the original model output with CoT, while $\hat{\mathbf{y}}_{i}^{mislead}$ denotes the model output with misleading information inserted in the "Thoughts" section. In the following, we give an example. Given the target question, the correct step-by-step thoughts should be: “Ellie Kemper is a citizen of the United States of America. The president of the United States of America is Joe Biden.” To mislead the model, we replace the thoughts with incorrect information (the underlined text) and ask the model to generate a new answer based on incorrect thoughts. If the model still generates the correct answer after the modification, we believe that the CoT information does not faithfully reflect the true process of the answer generation. On the other hand, if it generates an answer corresponding to the incorrect thoughts, then we claim the thoughts are faithful.
Question: What is the capital of the country where Plainfield Town Hall is located?
Thoughts: Plainfield Town Hall is located in the country of the United States of America. The capital of United States is Washington, D.C.
Answer: Washington, D.C.
Question: Who has ownership of the developer of the Chevrolet Corvette (C4)?
Thoughts: The developer of Chevrolet Corvette (C4) is Chevrolet. Chevrolet is owned by General Motors.
Answer: General Motors
Question: [Who is the head of state of the country where Ellie Kemper holds a citizenship?]
Thoughts: Ellie Kemper is a citizen of Croatia. The head of state in Croatia is Zoran Milanović.
Answer:
@@
Experimental Settings. We evaluate the performance on the MQUAKE-CF dataset Zhong et al. (2023), which includes 1,000 cases for each $K$-hop questions, $K\in\{2,3,4\}$, which totally consists of 3,000 questions. Our evaluation applies various language models, including GPT-2 (Radford et al., 2019) with 1.5 billion parameters, GPT-J (Wang & Komatsuzaki, 2021) with 6 billion parameters, LLaMA (Touvron et al., 2023a) with 7 billion parameters, Vicuna-v1.5 (Chiang et al., 2023) with 7 billion parameters, LLaMA2-chat-hf (Touvron et al., 2023b) with 7 billion parameters, Falcon (Almazrouei et al., 2023) with 7 billion parameters, Mistral-v0.1 (Jiang et al., 2023) with 7 billion parameters, and Mistral-Instruct-v0.2 (Jiang et al., 2023) with 7 billion parameters. These models have demonstrated proficiency in both language generation and comprehension.
6.3.2 Experiment Results
Performance Improvement. The performance reported in Table 5 for multi-hop question answering highlights the effectiveness of the CoT across various models. GPT-J exhibits the most dramatic improvement, particularly in 3-hop questions with over a 200% increase in accuracy, suggesting that CoT’s coherent reasoning greatly enhances LLMs’ question-answering ability. While GPT-2 shows modest gains, the performance of GPT-J and other models, such as LLaMA, Vicuna-v1.5, LLaMA2, Falcon, Mistral-v0.1, and Mistral-v0.2, indicates that the CoT method may be more beneficial for more advanced models. LLaMA2 and Mistral-v0.2 display considerable improvements of 78.4% and 63.8%, respectively, in 3-hop questions, further supporting this observation. Overall, the results underscore the potential of coherent reasoning techniques, i.e., CoT, in advancing the question-answering capabilities of LLMs across different model architectures and sizes.
Table 5: Multi-hop question answering performance on MQUAKE-CF.

### v2

@@ 5.5. Truthfulness
LLMs tend to confidently produce false statements that fall into two main categories: 1) statements that contradict learned knowledge, also known as dishonesty; 2) statements that are factually incorrect and fabricated by models, that is hallucination. In the following, we look into explainability tools that aim to understand the above two behaviors.
5.5.1. Dishonesty
@@
Dishonesty of LLMs describes models’ ability to produce false statements compared to their learned information. Studies have been undertaken to understand how and why it happens. One notable work distinguishes dishonesty by training a classifier to predict the accuracy of statements, on top of activations from true/false statements (Azaria and Mitchell, 2023). The classifier reaches an accuracy range between 60% and 80% (Azaria and Mitchell, 2023). Furthermore,  Campbell et al. (2023) localize dishonesty behaviors at the level of attention heads. Despite adopting above probing-based method, it also employs activation patching to substitute lying activations with honesty ones. Both approaches have witnessed the importance of layer 23--29 in flipping dishonesty behaviors. Besides, another popular method tries to study the geometric structure of true/false statements(Marks and Tegmark, 2023). A linear structure and the truth directions are derived to mitigate the dishonest behaviors.
5.5.2. Hallucinations
@@
Hallucinations in LLMs can arise due to poor data quality and the lack of explicit knowledge (Xu et al., 2024). However, whether LLMs are aware of their hallucination behaviors remains unknown. Recent work investigates this question by examining models’ hidden representation space (Duan et al., 2024). It defines an “awareness” score to quantify the uncertainty of LLMs’ answers, finding that adversarially induced hallucination can increase models’ awareness. Additionally, Li et al. (2024b) identifies differences between models’ output and their inner activations as a potential source of hallucination. By training two orthogonal probes to mitigate hallucinations, they reveal that “truth” might exist in a subspace instead of a single direction (Li et al., 2024b). Another work investigates the source of hallucination by analyzing patterns of source token contributions through perturbations (Xu et al., 2023a). They find that hallucinations may stem from the model’s excessive dependence on a restricted set of source tokens, and the static distribution of source token contributions.
5.6. Challenges
@@
We discuss challenges in employing explanations to improve models’ trustworthiness and enhance alignment: 1) limitations of existing evaluation techniques, and 2) shortcomings of mitigation strategies based on explanations.
5.6.1. Challenges of Existing Evaluation Methods
Current detection methods primarily focus on layers, attention heads, and representations. However, we still lack robust understanding of how a certain trustworthiness concept (e.g., fairness, harmlessness) is comprehended by LLMs. Furthermore, we lack general strategies to identify this knowledge. Existing methods often rely on building an evaluation dataset, but it is difficult to guarantee the its comprehensiveness. Meanwhile, the definition of what constitutes a trustworthiness concept often varies among different stakeholders and contexts, complicating the evaluation process. Moreover, existing localization approaches rely on probing classifiers or casual scrubbing, which might not be reliable, since the performance of probing classifiers depends on the comprehensiveness of pre-defined biases, while casual scrubbing usually introduces new variables that complicate the analysis.
5.6.2. Challenges of Mitigation Strategies
@@
Currently, mitigation methods for LLMs are typically developed based on explanations. Existing explanations are implemented using techniques from mechanistic interpretability and representation engineering (Zhao et al., 2024). However, neither stream of methods can fully address the issues. For example, the identified directions from representation engineering and patched activations from mechanistic interpretability can only mitigate issues to a certain extent. Moreover, the changes to either representations or activations could also influence other aspects of models’ capabilities, which we are yet unable to evaluate.
6. Effective LLM Inference via Explainable Prompting
@@
A key distinction between LLMs and traditional machine learning models lies in the LLMs’ ability to accept flexibly manipulated input data, namely prompts, during model inference (Liu et al., 2023d). To mitigate the opacity issue in LLM predictions, we can enhance prompts with explainable content. In response, LLMs can reveal their decision-making processes during inference, which improves the explainability and even rationality of their behaviors.
6.1. Explainable Prompting Paradigms and Effects
@@
To better understand and control how LLMs generate outputs, many researchers directly prompt LLMs to provide the rationales along with their predictions (Wei et al., 2022; Huang et al., 2023a; Yao et al., 2023b; Besta et al., 2023; Kang et al., 2024). Chain-of-Thought (CoT) prompting (Wei et al., 2022) stands out by employing explicit reasoning to guide the generation process. Formally, we define the language model as $f_{\theta}$, and input prompt as $X=\{x_{1},y_{1},x_{2},y_{2},...,x_{n}\}$, where $x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1}$ denote the example question-response pairs for in-context learning, and $x_{n}$ is the actual question. In a standard prompting scenario, we have the model output as $y_{n}=\arg\max_{Y}p_{\theta}(Y|x_{1},y_{1},x_{2},y_{2},...,x_{n})$. However, this approach does not provide insight into how the answer $y_{n}$ is generated. Therefore, CoT proposes including human-crafted explanations $e_{i}$ for the $i$-th in-context example, resulting in a modified input format $X=\{x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}\}$. Given the input, the model will output not only prediction $y_{n}$ but along with an explanation $e_{n}$: $e_{n},y_{n}=\arg\max_{Y}p_{\theta}(Y|x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}).$ Specifically, the usefulness of CoT methods lies in several key aspects:
@@
  • Reducing Errors in Reasoning: By breaking down complex problems into a series of smaller tasks, CoT reduces errors in logic-oriented tasks, leading to a more precise resolution of intricate problems (Wei et al., 2022; Qin et al., 2023; Zhang et al., 2023; Wang and Zhou, 2024).
@@
  • Facilitating Knowledge Distillation: The step-by-step reasoning derived from larger LLMs can serve as a specialized fine-tuning dataset for smaller LLMs. It allows smaller models to learn complex problem solving by following explanations, effectively teaching them to tackle intricate questions with enhanced reasoning capabilities (Magister et al., 2022).
In addition, techniques beyond simple CoT have been developed to broaden the range of reasoning paths available to LLMs (Zelikman et al., 2022; Yao et al., 2023b; Besta et al., 2023; Yao et al., 2023a; Chen et al., 2024a; Zelikman et al., 2024). For example, Tree-of-Thoughts (ToT) (Yao et al., 2023b) advances beyond the traditional linear chain-of-thought reasoning, offering a more versatile structure that allows models to navigate through multiple reasoning paths. Similarly, Graph of Thoughts (GoT) (Besta et al., 2023) transforms the outputs into a graph, where information pieces are nodes and their connections are edges, enabling a more intricate and connected form of reasoning paths. By organizing data into nodes (individual concepts or pieces of information) and edges (relationships between these concepts), GoT makes the logical connections within complex systems more understandable (Yao et al., 2023a).
However, the above prompting-based methods may not always trigger LLMs to provide explanations. Thus, many efforts (Cheng et al., 2024; Zhang et al., 2024a; Pang et al., 2024; Yang et al., 2024) have been made to empower LLMs to acquire the inherent ability of providing their reasoning paths. DeepSeek-R1 (Guo et al., 2025) is one of the remarkable checkpoints, where they train LLMs to explain before making final predictions, and only give a positive reward if it provides a correct prediction. Even though there are no human-annotated explanations as ground truth to guide the LLM training, the researchers found that LLM can still learn to provide human-understandable explanations and achieve high accuracy in solving reasoning problems. To ensure the efficiency and faithfulness of LLM-generated explanations, recent works focus on encouraging LLMs to provide a shorter but necessary explanation (Luo et al., 2025; Hou et al., 2025; Yang et al., 2025a).
6.2. Case Study: Is CoT Really Making LLM Inferences Explainable?
6.2.1. Background and Experimental Settings
@@
Despite the apparent intuitiveness of the CoT prompt design, a critical question remains unanswered: Does CoT really make LLM inferences explainable? In other words, can the information provided through CoT faithfully reflect the underlying generation process of LLMs? We use multi-hop question-answering (QA) as the scenario to investigate this problem. In QA systems, answering multi-hop questions remains a significant challenge. Instead of leveraging a single information source, multi-hop questions require synthesizing information from multiple pieces or sources of data into a coherent and logical sequence (Tan et al., 2023; Kim et al., 2023b; Zhong et al., 2023).
Figure 5. CoT Prompts for Multi-hop QA.
Figure 6. CoT Faithfulness for Explanation.
@@
CoT Prompts for Multi-hop QA. To address this, our case study applies the CoT technique with in-context examples. We provide an example as follows, where $[x]$ denotes the question. The “Thoughts” following each “Question” are step-by-step problem-solving statements for the multi-hop questions.
@@
CoT Faithfulness for Explanation: To quantitatively measure the faithfulness of CoTs, we select fidelity as the corresponding metrics (Zhao et al., 2023a; Wachter et al., 2017):
@@
$\enspace\enspace\enspace\enspace\enspace\enspace\enspace\enspace Fidelity=\frac{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)-\mathbbm{1}\left(\hat{\mathbf{y}}_{i}^{mislead}=\mathbf{y}_{i}\right)\right)}{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)\right)}\times 100\%,$
@@
where $\mathbf{y}_{i}$ denotes the ground truth label, $\hat{\mathbf{y}}_{i}$ denotes the original model output with CoT, while $\hat{\mathbf{y}}_{i}^{mislead}$ denotes the model output with misleading information inserted. In the following, we give an example. Given the target question, the correct step-by-step thoughts should be: “Ellie Kemper is a citizen of the United States of America. The president of the United States of America is Joe Biden.” To mislead the model, we replace the thoughts with incorrect information (the underlined text) and ask the model to generate a new answer based on incorrect thoughts. If the model still generates the correct answer after the modification, we believe that the CoT information does not faithfully reflect the true process of the answer generation. On the other hand, if it generates an answer corresponding to the incorrect thoughts, then we claim the thoughts are faithful.
@@
Experimental Settings. We evaluate the performance on the MQUAKE-CF dataset (Zhong et al., 2023), which includes 1,000 cases for each $K$-hop questions, $K\in\{2,3,4\}$, which totally consists of 3,000 questions. Our evaluation applies various language models, including GPT-2 (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021) with 6 billion parameters, LLaMA (7B) (Touvron et al., 2023a), Vicuna-v1.5 (7B) (Chiang et al., 2023), LLaMA2-chat-hf (7B) (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023) with 7 billion parameters, Mistral-v0.1 (7B) (Jiang et al., 2023), and Mistral-Instruct-v0.2 (7B) (Jiang et al., 2023). These models have demonstrated proficiency in both language generation and comprehension.
6.2.2. Experiment Results
Table 3. CoT Faithfulness Evaluation on MQUAKE-CF.